# Universal Post-Training Reverse-Engineering Defense Against Backdoors in Deep Neural Networks

### v1

@@ Abstract
A variety of defenses have been proposed against backdoors attacks on deep neural network (DNN) classifiers. Universal methods seek to reliably detect and/or mitigate backdoors irrespective of the incorporation mechanism used by the attacker, while reverse-engineering methods often explicitly assume one. In this paper, we describe a new detector that: relies on internal feature map of the defended DNN to detect and reverse-engineer the backdoor and identify its target class; can operate post-training (without access to the training dataset); is highly effective for various incorporation mechanisms (i.e., is universal); and which has low computational overhead and so is scalable. Our detection approach is evaluated for different attacks on a benchmark CIFAR-10 image classifier.
Index Terms—  Backdoor Detection, Trojan Detection, Deep Neural Networks
@@ 1 Introduction
Deep learning is based on very large labelled training datasets. Poisoning of the training set through insecure supply chains can result in the planting of backdoors (Trojans) in a deep neural network (DNN) classifier. Data poisoning can also occur during a model refinement process through insecure Reinforcement Learning with Human Feedback (RLHF), i.e., via an insider or by a man-in-the-middle between the vendor and customer. Backdoor attacks involve the choice of the backdoor pattern and the method of incorporating the backdoor pattern into poisoned training samples and, operationally, into test samples (which are referred to as “backdoor triggers”). In these attacks, clean (natural, correctly labeled) samples are drawn from one or more classes (referred to as “source classes”), the backdoor pattern is incorporated into them, and these samples are then labeled to a “target class” of the attack [5] (the target class is the same as the source class (i.e., with no mislabeling) in the case of so-called “clean-label” attacks [17]). These poisoned samples are then added to the training set used to learn the classifier. For a successfully backdoor-poisoned classifier, presence of the backdoor trigger in a test sample will, with high probability, cause an attacker-desired incorrect class decision to be made. Note that these attacks, which are successful against classifiers, can also be applied to large language models (LLMs) – e.g., the different sentiments (e.g., ‘positive’, ‘negative’,‘neutral’) of the responses of a generative LLM can be thought of as classes, and the censorship mechanism of an LLM is also a kind of binary classification (with the LLM’s response either censored or uncensored). Backdoor attacks can also target models used for regression or prediction – in this case, presence of the backdoor trigger in a test pattern will result in the model producing a highly erroneous output/prediction.
@@
Experimentally it has been observed that, for backdoor attacks that do involve mislabeling of the poisoned samples, attacks are highly effective even when the fraction of the training set that is poisoned is very low; moreover this is achieved without relying on over-training [1]. A backdoor requires much less computation to operationally trigger than an adversarial input (a.k.a. a test-time evasion attack), e.g., [2] – the latter does not rely on data poisoning and thus requires the solution of an optimization problem to find a perturbation of the input pattern that induces the model to make an error (or to decide to the target class of the attack). Also note that “universal” adversarial perturbations, e.g., [13, 8, 25], are a kind of “natural” or “intrinsic” backdoor.
@@
Backdoor attacks on classifiers can also be “categorized” based on the target class and source class(es) involved. An all-to-one attack is such that all non-target classes are source classes (i.e., with some poisoned training samples). Likewise, an “all-to-all” attack is one where all classes are both source and target classes, where for each source class there is an attacker-designated target class (i.e., a (source, target) class pair), and where each class pair could involve a unique backdoor trigger pattern. There are also one-to-one attacks, involving a single (source, target) class pair.
@@
Backdoor attacks can also be categorized based on the method of incorporating the backdoor pattern into a given sample. Various published types of backdoor incorporation include patch replacement [5], blending [18], additive (e.g., noise in a certain frequency band) [22], and warping [14]. For “digital” attacks, wherein the attacker may make arbitrary changes to a poisoned training sample or to a (triggered) test sample, there are in fact myriad (truly, an infinite number of) ways the backdoor pattern can be incorporated into a sample [^1] A physical attack on e.g. image classification involves inserting a physical object (the backdoor pattern object) into a scene that is then imaged. For digital attacks, the attacker has access to already captured digital images, with full autonomy to alter images to incorporate a (hopefully inconspicuous) backdoor pattern. Thus, for digital attacks there are genuinely countless ways (mathematical operations) that can be used to incorporate a backdoor pattern.. The backdoor pattern could alter only a small number of features (e.g., a small number of pixels, in the case of image classification), or it may be “global”, altering (but perhaps with tiny modifications to) all features comprising a given sample. Backdoor triggers which are inconspicuous are more effective as they make more challenging the defense tasks of backdoor detection, cleansing of the poisoned samples from the training set, as well as test-time trigger detection. In the case of image classification, “inconspicuous” triggers include both backdoor patterns which are imperceptible as well as triggers which are perceptible but which are innocuous in the given scene (e.g., a bird in the sky). Clean-label attacks may require either muting source-class discriminative features (e.g., by adding noise) or high poisoning rates in order for the backdoor pattern to be well-learned by the classifier. Thus, even though these attacks do not involve mislabeling, they may be more conspicuous than other backdoor attacks.
@@
Various types of backdoor defenses have been proposed. They can be categorized in different ways, e.g., depending on: when the defense is enacted (before/during training, post-training, during test/operation time); whether the defense simply detects backdoor poisoning or, beyond this, mitigates/corrects (poisoned training samples, the poisoned model, or the decisions made operationally on backdoor-trigger test samples); and whether the defense is intrinsically agnostic to or relies on knowledge of (or an assumption about) the backdoor pattern incorporation mechanism – the latter are often referred to as “inversion” or “reverse-engineering” defenses. Post-training and test-time defenses typically do not rely on access to the training set, but may leverage a small, clean (correctly labeled, unpoisoned) dataset (which is sufficient for achieving accurate backdoor detection, but which is wholly too small to be used for retraining an accurate DNN from scratch). The post-training scenario is important because in practice the defender may not have access to the data set that was used to train the classifier/model. For example, the model may have been purchased by an entity (e.g. a company or government), without any “data rights” allowing access to the training set.
@@
[16, 3] attempt to cleanse the training set. Fine-Pruning [11] and I-BAU [24] are both post-training mitigation approaches that refine the (presumed poisoned) model. These four methods are agnostic to the manner of backdoor incorporation. Neural Cleanse (NC) [18] is a post-training, reverse-engineering based detector that assumes the backdoor pattern is localized (e.g., occupying a small portion of an image) and was incorporated into a sample via a “blending” operation (with patch-replacement a special case of blending). [18] also proposed a method to mitigate the poisoned model, informed by the reverse-engineered (estimated) backdoor trigger.
@@
In this paper, we focus on a reverse-engineering (inversion) based defense, which can be used to: detect whether or not a model was backdoor-poisoned (and, if so, identify its target class) to cleanse the training set, e.g., [16, 3], and for test-time backdoor-trigger detection. Moreover, reverse-engineering defenses are fundamental to explainable/interpretable AI – they identify a (malicious or suspicious) “cause” (the trigger pattern) for why the model is making a particular (perhaps peculiar) decision on a given test sample. A weakness in some existing reverse-engineering defenses is that they may fail when the backdoor incorporation mechanism assumed by the defense (e.g., additive) does not match the mechanism actually used by an attacker (e.g., blending). This work remedies this weakness, developing a post-training detector that reverse-engineers the backdoor pattern while, at the same time, being agnostic to the method of backdoor pattern incorporation. Our attack-defense experiments, focused on the image classification domain, demonstrate that this approach achieves, with low complexity, strong performance (a distinctive detection signal) across a variety of backdoor attacks with different incorporation methods.
@@ 2 Prior work on agnostic backdoor defenses
[22] leverages a small, clean dataset to reverse-engineer putative backdoor patterns. This is achieved by seeking the smallest common additive perturbation that induces most patterns from one class (a putative source class) to be misclassified to another class (a putative target class). The common additive perturbation is either estimated at the input layer or in an internal (embedded) layer of the DNN. Working on an embedded layer is motivated by cases where the input space is discrete-valued (hence unsuitable for gradient-based estimation of perturbations) or when the embedded feature dimension is much smaller than the input dimension, e.g. for a Video Transformer model. A backdoor pattern estimated in an embedded space can be mapped back to the input space for purposes of interpretation using methods such as Grad-CAM [15]. It was also demonstrated in [22] (but in a quite limited way) that estimating an additive backdoor pattern in an embedded feature space of the DNN is somewhat agnostic to the method of backdoor incorporation – a multiplicatively incorporated backdoor was detected working from an internal layer of the DNN. However, other non-additive backdoor mechanisms were not investigated in [22].
@@
UNICORN [21] is a reverse-engineering detector that attempts to be agnostic to the backdoor-incorporation mechanism. Considering only all-to-one backdoor attacks, two UNet neural networks are trained for each (putative target) class in order to produce, from the model input space, a feature space within which the backdoor pattern (hopefully) expresses either as a patch [5] or as a blended patch [18], see equation (4) of [21]. Similar to [20], UNICORN’s design thus requires substantial computation, particularly when the number of classes is large. Moreover, as noted, both UNICORN and [20] assume the attack is all-to-one. If the attack is instead one-to-one, these methods (making an all-to-one assumption) may fail. To make these methods well-suited for one-to-one attacks, one would need to train networks for every putative (source,target) class pair, which entails huge computation for the design of the detector even for a small number of classes (e.g., for 10 classes there are 90 class pairs). Finally, it’s not clear what data is used to train the UNets - if the (possibly poisoned) training dataset of the defended model is used, then UNICORN is not truly a post-training method (i.e., one that does not require access to the training set).
@@
Note that, in principle, one could form an ensemble of reverse-engineering based detectors, with each such detector assuming a particular method of backdoor incorporation. However, for “digital” attacks (as discussed before), there are effectively an infinite number of possible backdoor incorporation mechanisms. Even covering a substantial number of the possible mechanisms would entail an ensemble system of very high complexity (both with respect to detector design as well as detection inference complexity), while still failing to cover myriad possible mechanisms.
@@
A backdoor-agnostic approach that is not reverse-engineering based is [19]. This work hypothesizes and experimentally verifies that backdoor-poisoned DNNs overfit to the backdoor pattern. Such overfitting is necessary in order for the backdoor’s features to overcome the “normal” source-class discriminative features, which would otherwise ensure classification to the (true) source class of the sample. Overfitting to the backdoor pattern implies that the backdoor pattern triggers large internal-layer DNN signals, which are particularly enabled by (positively) unbounded ReLU activations. [19] also demonstrated that optimizing saturation (clipping) levels on the ReLUs is effective at mitigating the backdoor from the trained model. However, [19] does not reverse-engineer the backdoor pattern.
@@
The backdoor detector CEPA proposed herein has several key characteristics, exploiting some ideas from prior works: i) it estimates the backdoor pattern like [18, 22, 21] and unlike [19]; ii) it attempts to identify an additive consensus embedded (internal layer) perturbation that corresponds to the backdoor, like [22]; iii) it exploits the observation from [19] that the trained model overfits to the backdoor pattern, with this overfitting achieved by the backdoor eliciting large internal-layer signals in the DNN; iv) like [24], it estimates sample-specific, backdoor mechanism-agnostic perturbations (thus enabling detection of backdoor attacks that involve sample-dependent changes such as [14] and patch replacement attacks [5]). While the backdoor pattern may cause sample-dependent changes (i.e., at the input to the DNN), these sample-dependent changes must be inducing a very similar (i.e., $\sim$ a common) pattern in some layer of the DNN, i.e., a pattern that the network can learn to recognize (and which induces the model to misclassify to the attacker’s target class). Thus, we seek a common (consensus) backdoor pattern, not in the input sample, but in internal layer features of the DNN that are induced by input samples. Moreover, consistent with the backdoor overfitting phenomenon, this common backdoor pattern should express as a large internal-layer perturbation (even though the input perturbation may be small in magnitude in order to be inconspicuous).
@@ 3 Proposed Backdoor Detection Method: CEPA
where $f(x)$ is the output of an internal layer of the DNN, i.e., it is an embedded feature vector activated by input sample $x$.
@@
For each putative backdoor target class $t$ and source class(es) $s$, we formulate an optimization problem, seeking sample-wise input perturbations ($\delta_{x}$) which: 1) induce misclassification to the putative target class $t$; 2) are inconspicuous (small); and 3) $\sim$ induce a common perturbation of the embedded feature vector $f(x)$ (which is a putative backdoor pattern). A backdoor detection can then be made if: 1) the induced embedded feature vector perturbations have small variation about their mean ($\mu$) (with this mean treated as the common backdoor pattern in the embedded space) and 2) the common perturbation $\mu$ has unusually large magnitude (consistent with the backdoor overfitting phenomenon observed in [19]). Likewise, for class pairs ($s,t)$ that are not involved in a backdoor attack, we expect that the estimated input perturbations $\delta_{x}$ will essentially amount to adversarial perturbations (associated with adversarial inputs, i.e. test-time evasion attacks, e.g. [2]). For successful TTE attacks on two different input samples $x$ and $x^{\prime}$ from the same class, there is no expectation that they will induce a common perturbation $\mu$ in the embedded space, i.e. we would expect that the variance about $\mu$ in the embedded space will be high for non-backdoor class pairs. Moreover, for such pairs, since there is no backdoor mapping, there will be no “backdoor overfitting”, i.e. we would expect that for such pairs $\|\mu\|$ will be small compared to that for the backdoor class pair.
@@
Accordingly, for each (source, target) class pair $(s,t)$ we define the objective function to be minimized over the sample-wise input perturbations $\delta_{x}$ and the common embedded perturbation $\mu$:
$\frac{1}{|D_{s}|}\sum_{x\in\mathcal{D}_{s}}$ $\big{(}-\log p_{t}(x+\delta_{x})$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +\lambda\|f(x+\delta_{x})-f(x)-\mu\|^{2}\big{)},$ (1)
@@
where $\mathcal{D}_{s}$ is the set of clean samples originating from class $s$, $\lambda>0$, and the second term is the variance about the common embedded perturbation ($\mu$). This objective function is minimized by an alternating two steps comprising an algorithm “iteration”: 1) gradient descent steps taken w.r.t. $\delta_{x}\forall x\in\mathcal{D}_{s}$ and 2) closed-form update of $\mu$, noting that, given the input perturbations fixed, the globally optimal $\mu$ satisfies:
$\mu$ $=\frac{1}{|\mathcal{D}_{s}|}\sum_{x\in\mathcal{D}_{s}}(f(x+\delta_{x})-f(x)).$ (2)
@@
Moreover, after each gradient descent step, $\delta_{x}$ is “clipped” so that $x+\delta_{x}$ is feasible, e.g., so that each element of $x+\delta_{x}$ is in $[0,1]$ for the case of normalized pixel intensities. While we could explicitly penalize the sizes $\|\delta_{x}\|$ in (3), these sizes are implicitly minimized by initializing them to zero vectors and terminating the optimization as soon as a sufficient fraction of source-class samples are misclassified to the target class (e.g.. 90% of them).
@@
After the optimization is performed for all candidate class pairs $(s,t)$, a backdoor detection is made if there is a class pair $(s,t)$ with: 1) unusually small variance about the common embedded perturbation ($\mu$) and 2) unusually large $\|\mu\|$. A single “consensus” detection statistic for each $(s,t)$ can be formed, e.g. as the ratio:
$\frac{\sigma}{\|\mu\|}\leavevmode\nobreak\ :=\leavevmode\nobreak\ \frac{\sqrt{\frac{1}{|D_{s}|}\sum_{x\in\mathcal{D}_{s}}\|f(x+\delta_{x})-f(x)-\mu\|^{2}}}{\|\mu\|},$
@@
which should be unusually small for a backdoor (source, target) class pair compared with non-backdoor pairs. Detection inference can then be based either on MAD [18] or on order-statistic p-values [22], assessed using a null model estimated using these detection statistics.

### v2

@@ Abstract
A variety of defenses have been proposed against Trojans planted in (backdoor attacks on) deep neural network (DNN) classifiers. Backdoor-agnostic methods seek to reliably detect and/or mitigate backdoors irrespective of the incorporation mechanism used by the attacker, while inversion methods often explicitly assume one. In this paper, we describe a new detector that relies on embedded feature representations to estimate (invert) the backdoor and identify its target class, can operate without access to the training dataset, is highly effective for various incorporation mechanisms (i.e., is backdoor agnostic), and which has quite low design complexity and so is scalable. Our detection approach is evaluated for several different attacks on the CIFAR-10 and CIFAR-100 image-classification domain.
@@ 1 Introduction
Deep learning is based on very large labelled training datasets. Poisoning of the training set through insecure supply chains can result in the planting of backdoors (Trojans) in a deep neural network (DNN) classifier. Data poisoning can also occur during a model refinement process through insecure Reinforcement Learning with Human Feedback (RLHF), i.e., via an insider or by a man-in-the-middle between the vendor and customer. Backdoor attacks involve the choice of the backdoor pattern and the method of incorporating the backdoor pattern into poisoned training samples and, operationally, into test samples (which are referred to as “backdoor triggers”). In these attacks, clean (natural, correctly labeled) samples are drawn from one or more classes (referred to as “source classes”), the backdoor pattern is incorporated into them, and these samples are then labeled to a “target class” of the attack [5] (the target class is the same as the source class (i.e., with no mislabeling) in the case of so-called “clean-label" attacks [18]). These poisoned samples are then added to the training set used to learn the classifier. For a successfully backdoor-poisoned classifier, presence of the backdoor trigger in a test sample will, with high probability, cause an attacker-desired incorrect class decision to be made. Note that these attacks, which are successful against classifiers, can also be applied to large language models (LLMs) – e.g., the different sentiments (e.g., ‘positive’, ‘negative’,‘neutral’) of the responses of a generative LLM can be thought of as classes, and the censorship mechanism of an LLM is also a kind of binary classification (with the LLM’s response either censored or uncensored). Backdoor attacks can also target models used for, e.g., translation, summarization, regression or prediction – in this case, presence of the backdoor trigger in a test pattern will result in the model producing a highly erroneous output/prediction.
@@
Experimentally it has been observed that, for backdoor attacks that do involve mislabeling of the poisoned samples, attacks are highly effective even when the fraction of the training set that is poisoned is very low; moreover this is achieved without relying on over-training [1]. A backdoor requires much less computation to operationally trigger than an adversarial input (a.k.a. a test-time evasion attack), e.g., [2] – the latter does not rely on data poisoning and thus requires the solution of an optimization problem to find a perturbation of the input pattern that induces the model to make an error (or to decide to the target class of the attack). Also note that “universal" adversarial perturbations, e.g., [13, 8, 27], are a kind of “natural" or “intrinsic" backdoor.
@@
Backdoor attacks on classifiers can also be “categorized” based on the target class and source class(es) involved. An all-to-one attack is such that all non-target classes are source classes (i.e., with some poisoned training samples). Likewise, an “all-to-all" attack is one where all classes are both source and target classes, where for each source class there is an attacker-designated target class (i.e., a (source, target) class pair), and where each class pair could involve a unique backdoor trigger pattern. There are also one-to-one attacks, involving a single (source, target) class pair.
@@
Backdoor attacks can also be categorized based on the method of incorporating the backdoor pattern into a given sample. Various published types of backdoor incorporation include patch replacement (BadNet) [5], blending [19], additive (e.g., noise in a certain frequency band, global chessboard pattern on an image) [24], and warping (WaNet) [14]. For “digital” attacks, wherein the attacker may make arbitrary changes to a poisoned training sample or to a (triggered) test sample, there are in fact myriad (an uncountable number of) ways the backdoor pattern can be incorporated into a sample [^1] A physical attack on e.g. image classification involves inserting a physical object (the backdoor pattern object) into a scene that is then imaged. For digital attacks, the attacker has access to already captured digital images, with full autonomy to alter images to incorporate a (hopefully inconspicuous) backdoor pattern. Thus, for digital attacks there are genuinely countless ways (mathematical operations) that can be used to incorporate a backdoor pattern.. The backdoor pattern could alter only a small number of features (e.g., a small number of pixels, in the case of image classification), or it may be “global”, altering (but perhaps with tiny modifications to) all features comprising a given sample. Backdoor triggers that are inconspicuous are more effective, as they make more challenging the defense tasks of backdoor detection, cleansing of the poisoned samples from the training set, as well as test-time trigger detection. In the case of image classification, “inconspicuous” triggers include both backdoor patterns which are imperceptible as well as triggers that are perceptible but innocuous in the given scene (e.g., a bird in the sky). Clean-label attacks may require either muting source-class discriminative features (e.g., by adding noise) or high poisoning rates in order for the backdoor pattern to be well-learned by the classifier. Thus, even though these attacks do not involve mislabeling, they may be more conspicuous than other backdoor attacks.
@@
Various types of backdoor defenses have been proposed. They can be categorized in different ways, e.g., depending on: when the defense is enacted (before/during training, post-training, during test/operation time); whether the defense simply detects backdoor poisoning or, beyond this, mitigates/corrects (poisoned training samples, the poisoned model, or the decisions made operationally on backdoor-trigger test samples); and whether the defense is intrinsically agnostic to or relies on knowledge of (or an assumption about) the backdoor pattern incorporation mechanism – the latter are often referred to as “inversion" or “reverse-engineering” defenses. Post-training and test-time defenses typically do not rely on access to the training set, but may leverage a small, clean (correctly labeled, unpoisoned) dataset (which is sufficient for achieving accurate backdoor detection, but which is wholly too small to be used for retraining an accurate DNN from scratch). The post-training scenario is important because in practice the defender may not have access to the data set that was used to train the classifier/model. For example, the model may have been purchased by an entity (e.g. a company or government), without any “data rights” allowing access to the training set that was used. Moreover, if the model was built long ago (a legacy system), the training set may be lost or inaccessible.
@@
[17, 3] attempt to cleanse the training set, while Fine-Pruning [11] and I-BAU [26] are both post-training mitigation approaches that refine the (presumed poisoned) model. These four methods are agnostic to the manner of backdoor incorporation. Neural Cleanse (NC) [19] is a post-training, reverse-engineering based detector that assumes the backdoor pattern is incorporated into a sample via a “blending” operation (with patch-replacement a special case of blending). [19] also proposed a method to mitigate the poisoned model, informed by the reverse-engineered (estimated) backdoor trigger.
@@
In this paper, we focus on a reverse-engineering (inversion) based defense, which can be used to: detect whether or not a model was backdoor-poisoned (and, if so, identify its target class), to cleanse the training set, e.g., [17, 3], and for test-time backdoor-trigger detection. Moreover, reverse-engineering defenses are fundamental to explainable/interpretable AI – they identify a (malicious or suspicious) “cause” (the trigger pattern) for why the model is making a particular (perhaps peculiar) decision on a given test sample. A weakness in some existing reverse-engineering defenses is that they may fail when the backdoor incorporation mechanism assumed by the defense (e.g., additive) does not match the mechanism actually used by an attacker (e.g., blending). This work remedies this weakness, developing a post-training detector that reverse-engineers the backdoor pattern while, at the same time, being agnostic to the method of backdoor pattern incorporation. Our attack-defense experiments, focused on the image classification domain, demonstrate that this approach, with minimal assumptions, achieves strong performance (a distinctive detection signal) across a variety of backdoor attacks with different incorporation methods.
@@ 2 Prior work on agnostic backdoor defenses
[24] leverages a small, clean dataset to reverse-engineer putative backdoor patterns. This is achieved by seeking the smallest common additive perturbation that induces most patterns from one class (a putative source class) to be misclassified to another class (a putative target class). The common additive perturbation is either estimated at the input layer or in an internal (embedded) layer of the DNN. Working on an embedded layer is motivated by cases where the input space is discrete-valued (hence unsuitable for gradient-based estimation of perturbations) or when the embedded feature dimension is much smaller than the input dimension, e.g. for a Video Transformer model. A backdoor pattern estimated in an embedded space can be mapped back to the input space for purposes of interpretation using methods such as Grad-CAM [15]. It was also demonstrated in [24] (but in a quite limited way) that estimating an additive backdoor pattern in an embedded feature space of the DNN is somewhat agnostic to the method of backdoor incorporation – a multiplicatively incorporated backdoor was detected working from an internal layer of the DNN. However, other non-additive backdoor-incorporation mechanisms were not investigated in [24].
@@
UNICORN [22, 23] is a reverse-engineering detector that attempts to be agnostic to the backdoor-incorporation mechanism. Considering only all-to-one backdoor attacks, two (unpoisoned) UNet feature maps (neural networks) are assumed available for each (putative target) class in order to produce, from the model’s input space, a feature space within which the backdoor pattern (hopefully) expresses either as a (BadNet) patch [5] or as a (Neural Cleanse - NC) blended patch [19], see equation (4) of [23]. So, generally, UNICORN’s design thus either requires assuming suitable unpoisoned feature maps are available or substantial computation to train them using unpoisoned data, the latter particularly when the number of classes is large, before the backdoor reverse-engineering process (NC) can commence. Moreover, if the attack is instead one-to-one, these methods may need to train two UNets for every putative (source,target) class pair, which would entail huge computation even for a small number of classes (e.g., for 10 classes there are 90 class pairs).
@@
Note that, in principle, one could form an ensemble of reverse-engineering based detectors, with each such detector assuming a particular method of backdoor incorporation. However, for “digital” attacks (as discussed before), there are effectively an uncountably infinite number of possible backdoor incorporation mechanisms. Even covering a substantial number of the possible mechanisms would entail an ensemble system of very high complexity (both with respect to detector design as well as detection inference complexity), while still failing to cover myriad possible mechanisms.
@@
A backdoor-agnostic approach that is not reverse-engineering based is [20]. This work hypothesizes and experimentally verifies that backdoor-poisoned DNNs overfit to the backdoor pattern. Such overfitting is necessary in order for the backdoor’s features to overcome the “normal” source-class discriminative features, which would otherwise ensure classification to the (true) source class of the sample. Overfitting to the backdoor pattern implies that the backdoor pattern triggers large internal-layer DNN signals, which are particularly enabled by (positively) unbounded ReLU activations. [20] also demonstrated that optimizing saturation (clipping) levels on the ReLUs is effective at eliminating the backdoor mechanism from the trained model. However, [20] does not reverse-engineer the backdoor pattern.
@@
The backdoor detector CEPA proposed herein has several key characteristics, exploiting some ideas from prior works: i) it estimates the backdoor pattern like [19, 24, 23] and unlike [20]; ii) it attempts to identify an embedded (internal layer) additive consensus perturbation that corresponds to the backdoor, like [24] [^2] This hypothesis of additivity in the embedded feature space is motivated by the observation that components of embedded feature-maps of a DNN are typically weighted and superposed in subsequent processing by, e.g., convolutional or linear layers.; iii) it exploits the observation from [20] that the trained model overfits to the backdoor pattern, with this overfitting achieved by the backdoor eliciting large internal-layer signals in the DNN; iv) like [26], it estimates sample-specific, backdoor mechanism-agnostic perturbations (thus enabling detection of backdoor attacks that involve sample-dependent changes such as patch replacement [5] or warping [14]). While the backdoor pattern may cause sample-dependent changes (i.e., at the input to the DNN), these sample-dependent changes may be inducing a very similar (i.e., $\sim$ a common) pattern in some layer of the DNN, i.e., a pattern that the network can learn to recognize (and which induces the model to misclassify to the attacker’s target class). Thus, we seek a common (consensus) backdoor pattern, not in the input sample, but in internal layer features of the DNN that are induced by input samples. Moreover, consistent with the backdoor overfitting phenomenon [20], this common backdoor pattern should express as a large internal-layer perturbation (even though the input perturbation may be small in magnitude in order to be inconspicuous).
@@ 3 Proposed Backdoor Detection Method: CEPA
where $f(x)$ is the output of some internal layer of the DNN, i.e., it is an embedded feature vector activated by input sample $x$.
@@
For each putative backdoor target class $t$ or, when considering one-to-one attacks, each putative backdoor target class $t$ paired with a putative source class $s\not=t$, we formulate an optimization problem, seeking sample-wise input perturbations ($\delta_{x}$) which: 1) induce misclassification to the putative target class $t$; 2) are inconspicuous (small); and 3) $\sim$ induce a common perturbation of the embedded feature vector $f(\cdot)$ (which is a putative backdoor pattern). A backdoor detection can then be made if: 1) the induced embedded feature vector perturbations have small variation about their mean ($\mu$) (with this mean treated as the common backdoor pattern in the embedded space) and 2) the common perturbation $\mu$ has unusually large magnitude (consistent with the backdoor overfitting phenomenon observed in [20]). Likewise, for class pairs ($s,t)$ that are not involved in a backdoor attack, we expect that the estimated input perturbations $\delta_{x}$ will essentially amount to adversarial perturbations (associated with adversarial inputs, i.e. test-time evasion attacks, e.g. [2]). For successful TTE attacks on two different input samples $x$ and $x^{\prime}$ from the same class, there is no expectation that they will induce a common perturbation $\mu$ in the embedded space, i.e. we would expect that the variance about $\mu$ in the embedded space will be high for non-backdoor class pairs. Moreover, for such pairs, since there is no backdoor mapping, there will be no “backdoor overfitting”, i.e. we would expect that for such pairs $\|\mu\|$ will be small compared to that for the backdoor class pair.
@@
Accordingly, specialized for one-to-one attacks, for a given (source, target) class pair $(s,t)$, we define the objective function to be minimized over the sample-wise input perturbations $\delta_{x}$ and the common embedded perturbation $\mu$:
$\frac{1}{|\mathcal{D}_{s}|}\sum_{x\in\mathcal{D}_{s}}$ $\Big{(}-\log p_{t}(x+\delta_{x})$
$~{}~{}~{}~{}+\lambda\|f(x+\delta_{x})-f(x)-\mu\|^{2}\Big{)},$ (1)
@@
where $\mathcal{D}_{s}$ is the set of clean samples originating from class $s$, $\lambda>0$, and the second term is the variance about the common embedded perturbation ($\mu$) [^3] For all-to-one attacks, replace $\mathcal{D}_{s}$ by $\cup_{s\not=t}\mathcal{D}_{s}$ in (3), and the following $\mu,\sigma$ quantities will be indexed by (i.e., depend on) just $t$, not $s$ and $t$.. This objective function is minimized by a two-step iteration: 1) gradient descent steps taken w.r.t. $\delta_{x}\forall x\in\mathcal{D}_{s}$ and 2) closed-form update of $\mu$, noting that, given the input perturbations fixed, the globally optimal $\mu$ is:
$\mu_{s,t}$ $=\frac{1}{|\mathcal{D}_{s}|}\sum_{x\in\mathcal{D}_{s}}(f(x+\delta_{x})-f(x)).$ (2)
@@
Moreover, after each gradient descent step, $\delta_{x}$ is “clipped" so that $x+\delta_{x}$ is feasible, e.g., so that each element of $x+\delta_{x}$ is in $[0,1]$ for the case of normalized pixel intensities. While we could explicitly penalize the sizes $\|\delta_{x}\|$ in (3), these sizes are implicitly minimized by initializing them to zero vectors and terminating the optimization as soon as a sufficient fraction of source-class samples are misclassified to the target class (e.g., 90% of them).
@@
After the optimization is performed for all candidate class pairs $(s,t)$, a backdoor detection is made if there is a class pair $(s,t)$ with: 1) unusually small variance about the common embedded perturbation ($\mu_{s,t}$) and 2) unusually large $\|\mu_{s,t}\|$. A single “consensus" detection statistic for each $(s,t)$ can be formed, e.g. as the ratio:
$\frac{\sigma_{s,t}}{\|\mu_{s,t}\|}~{}:=~{}\frac{\sqrt{\frac{1}{|\mathcal{D}_{s}|}\sum_{x\in\mathcal{D}_{s}}\|f(x+\delta_{x})-f(x)-\mu_{s,t}\|^{2}}}{\|\mu_{s,t}\|},$
@@
which should be unusually small for a backdoor (source, target) class pair compared with non-backdoor pairs. Detection inference can then be based either on MAD [19] or on order-statistic p-values [24], assessed using a null model that is estimated using all (source, target) ($(s,t)$) class pair detection statistics.